OpenStudy (anonymous):

pleasee i need help asap with complex numbers.. compute (3i + i)/1-2i in the form of a+bi

OpenStudy (anonymous):

\[\frac{4i}{1-2i}*\frac{1+2i}{1+2i}\]

OpenStudy (anonymous):

\[\frac{-8+4i}{5}\]

OpenStudy (anonymous):

how did u get that ?

OpenStudy (anonymous):

in general multiply top and bottom by the conjugate of the bottom. the conjugate of \[a+bi\] is \[a-bi\] and this works because \[(a+bi)(a-bi)=a^2+b^2\] a real number

OpenStudy (anonymous):

so in your example \[(1+2i)(1-2i)=1^2+2^2=5\]

OpenStudy (anonymous):

all the real work is multiplying in the numerator

OpenStudy (anonymous):

what about the numeratorr?

OpenStudy (anonymous):

(-8/5)+4i/5